On January 1, Hemi Corporation issued $500,000 of 15-year, 10% bonds for $586,460 yielding an effective interest rate(yield to maturity) of 8%. Interest is paid semi-annually on June 30 and December 31/

Required:

1) Show computations to confirm the issue price.

2) Compute an amortization schedule for the 15 years of the bonds' life.

3) Give Journal entries for: (A) bond issuance, (b) semi-annual interest payment on June 30 of the first year, and (c) the semi-annual interest payment on December 31 of the first year.
Answer #1
Bond characterstics Amount
1) Principal 500,000
interest           25,000
Market interest rate 4%
periods to maturity 30
issue price 586,460
principal * PV of $1 at 4% for 30 yrs =
500,000 * 0.30832        = 154160
interest * PV of ordinary annuity at 4%=
25000 * 17.29203 = 432300
bond issue price 586460
2) Amortization Schedule
Date cash intt inttt exp premium Carrying
amortized value
1/1/ 586,460
30-Jun      25,000 23458 1542 584918
31-Dec      25,000 23397 1603 583315
30-Jun      25,000 23333 1667 581648
31-Dec      25,000 23266 1734 579914
30-Jun      25,000 23197 1803 578110
31-Dec      25,000 23124 1876 576235
30-Jun      25,000 23049 1951 574284
31-Dec      25,000 22971 2029 572255
30-Jun      25,000 22890 2110 570146
31-Dec      25,000 22806 2194 567951
30-Jun      25,000 22718 2282 565669
31-Dec      25,000 22627 2373 563296
30-Jun      25,000 22532 2468 560828
31-Dec      25,000 22433 2567 558261
30-Jun      25,000 22330 2670 555592
31-Dec      25,000 22224 2776 552815
30-Jun      25,000 22113 2887 549928
31-Dec      25,000 21997 3003 546925
30-Jun      25,000 21877 3123 543802
31-Dec      25,000 21752 3248 540554
30-Jun      25,000 21622 3378 537176
31-Dec      25,000 21487 3513 533663
30-Jun      25,000 21347 3653 530010
31-Dec      25,000 21200 3800 526210
30-Jun      25,000 21048 3952 522259
31-Dec      25,000 20890 4110 518149
30-Jun      25,000 20726 4274 513875
31-Dec      25,000 20555 4445 509430
30-Jun      25,000 20377 4623 504807
31-Dec      25,000 20193 4807 500000
3) Date Account titles & Explanations Debit Credit
1-Jan Cash 586,460
Premium on bonds payable 86,460
Bonds payable 500,000
30-Jun interest expense 23458
premium on bonds payable 1542
cash      25,000
31-Dec interest expense 23397
premium on bonds payable 1603
cash      25,000